OpenStudy (anonymous):

the graph of y=bx^2-cx crosses the x-axis at the point (4, 0). The gradient at the point is 1. The value of c is ?!?!?!? please help me!

OpenStudy (anonymous):

b=1 y=0 x=4 plug the values in the equation and get c

hartnn (hartnn):

to find the point where a graph crosses 'x' axis, we put y=0 so, put y=0 in y=bx^2-cx and what you get ?

hartnn (hartnn):

and since it crosses at 4,0, you put x=4. you'll get one equation in b,c.

OpenStudy (anonymous):

i will try thak you :)

hartnn (hartnn):

ok, and to get other equation, differentiate the equation y=bx^2-cx and put y'=gradient =1, and x=4. Now you have 2 equations , 2 unknowns, easy to solve :)

OpenStudy (anonymous):

but am i able to get variable free answer in this quetion ?

hartnn (hartnn):

means ? show how much you could do ?

OpenStudy (anonymous):

ahh dont worry! i got it lol

hartnn (hartnn):

okay :)

OpenStudy (anonymous):

thank you so much for your help!

hartnn (hartnn):

welcome ^_^